y-4=3/5(x-2)
How do i graph this, or solve it ? im so confused if you could help it would be greatly apreciated. ! please, and thankss .

Answer by mananth(16946)

y-4=3/5(x-2)
multiply equation by 5
5(y-4)=3(x-2)
5y-20=3x-6
3x-5y=-14
3x+14=5y
generate values for y for different values of x
x= 0, y=14/5 (0,14/5)
x=14/3 , y =0 (-14/3, 0)


m.ananth@hotmail.ca